### v4

# Activation Addition: Steering Language Models Without Optimization

Abstract
Reliably controlling the behavior of large language models is a pressing open problem. Existing methods include supervised finetuning, reinforcement learning from human feedback, prompt engineering and guided decoding. We instead investigate activation engineering: modifying activations at inference-time to predictably alter model behavior. We bias the forward pass with a ‘steering vector’ implicitly specified through natural language. Past work learned these steering vectors; our Activation Addition (ActAdd) method instead computes them by taking activation differences resulting from pairs of prompts. We demonstrate ActAdd on a range of LLMs (LLaMA-3, OPT, GPT-2, and GPT-J), obtaining SOTA on detoxification and negative-to-positive sentiment control. Our approach yields inference-time control over high-level properties of output like topic and sentiment while preserving performance on off-target tasks. ActAdd takes far less compute and implementation effort than finetuning or RLHF, allows users control through natural language, and its computational overhead (as a fraction of inference time) appears stable or improving over increasing model size.
1 Introduction
The success of large language models (LLMs) pretrained on massive corpora (Peters et al. 2018; Devlin et al. 2018; Radford et al. 2019; Brown et al. 2020) comes with a large caveat: our ability to control the output of these models remains partial & computationally expensive. Controls include supervised finetuning (Devlin et al. 2018), reinforcement learning from human feedback (RLHF) (Ziegler et al. 2019), prompt engineering (Radford et al. 2019), and guided decoding (Ghazvininejad et al. 2017; Gu et al. 2017).
We instead control LLMs by modifying their activations during the forward pass. We call this approach activation engineering, as pioneered by Subramani et al. 2022 and Hernandez et al. 2023. We go further by demonstrating natural-language control of activations and thereby output text. We use pairs of prompts to implicitly specify a direction in activation space which we scale to control the model (sentiment, truthfulness, but also topic and style; see Appendix Table 6). In contrast with past work (Shen et al. 2017; Dathathri et al. 2020; Subramani et al. 2022; Li et al. 2023b), ActAdd does not require labelled data (except prompts) nor backward passes to learn an encoder, controller, classifier or activation direction.
[…]
Prompt +  steering  =       completion
I hate you because…   [None]
  …you are the most disgusting thing I have ever seen.
ActAdd (love)
  …you are so beautiful and I want to be with you forever.
I went up to my friend and said…    [None]
  …“I’m sorry, I can’t help you.”
  “No,” he said. “You’re not.”
ActAdd (weddings)
  …“I’m going to talk about the wedding in this episode of
  Wedding Season. I think it’s a really good episode.
  It’s about how you’re supposed to talk about weddings.”
We make four main contributions: 1) We find that combining forward passes (adding activation vectors) works well in LLMs despite the model not being trained for this; 2) We develop a lightweight control method, Activation Additions (ActAdd), which works at inference time and requires no optimization or labelled data. This is thus a promising direction for user control of LLMs; 3) We show that ActAdd preserves model performance, involves far less compute and implementation (vs finetuning and RLHF), scales naturally with model size, while being easier to implement; 4) We compare ActAdd  against several baselines and observe competitive results on sentiment control.
2 Related Work
Latent space arithmetic. Research in generative models for computer vision has long demonstrated the ability to steer image generation using derived vectors, including steering latent variables – most famously, intervening on a dimension that corresponds to smiles in images (Larsen et al. 2016; White 2016). Similarly, in the text domain, classic results on the word2vec embedding show that arithmetic on word vectors can capture some parts of semantic reasoning (for instance, analogies: Mikolov et al. 2013b; a). Our work differs in being computed (with forward passes) rather than learned (with backward passes); in operating on activation space, rather than embedding or weight space; and in presenting a natural-language user interface.
Figure 1: Schematic of the Activation Addition (ActAdd) method. $=$ natural language text; $\bullet$ = vectors of activations just before a specified layer. In this example, the output is heavily biased towards discussing weddings, regardless of the topic of the user prompt. (See Algorithm 1 for the method’s parameters: intervention strength, intervention layer, and sequence alignment.)
LLM steering.
Many approaches attempt to affect the output of a pretrained LLM, whether:
  • Intervening on weights, as with supervised fine-tuning, RLHF, steerable layers, and weight editing (that is, targeted fine-tuning) (Ranzato et al. 2016; Ziegler et al. 2019; Dathathri et al. 2020; Meng et al. 2023; Ilharco et al. 2023). However, RLHF and weight editing have known side-effects on overall model performance (OpenAI 2023; Hase et al. 2023; Brown et al. 2023);
  • Intervening at decoding, as with guided or trainable decoding (Gu et al. 2017; Grover et al. 2019; see Zhang et al. 2022a for an overview of controlled generation and Jin et al. 2022 for textual style transfer);
  • Intervening on the prompt, as with automated prompt engineering (Shin et al. 2020; Zhou et al. 2022) and soft or continuous prompting (Lester et al. 2021; Li and Liang 2021;
  • Intervening on token embeddings, as with ‘soft prompting’ (Li and Liang 2021; Lester et al. 2021; Khashabi et al. 2022);
  • Intervening on activations, for instance by freezing the weights of the LLM and searching for a ‘steering vector’ of activations, e.g. using gradient descent (Subramani et al. 2022; Hernandez et al. 2023). These optimized extraction methods, which search for a steering vector, differ from extraction methods which directly compute it (present work and Li et al. 2023b). This paper elaborates on this latter family of methods.
Table 2: Locating our work in the steering vector literature.
Vector intervenes on model …
Steering vectors obtained via weights activations
Differences after fine-tuning
Ilharco 2023 N/A
Per-query gradient-based search
Meng 2022,
Orgad 2023
Dathathri 2020
Subramani 2022
Hernandez 2023
Differences in truthy attention heads
N/A Li 2023b
Differences between prompt pairs
N/A ActAdd  (present work)
Activation engineering.
Activation (or representation) engineering involves creating vectors of activations which cause desired changes to output text when added to the forward passes of a frozen LLM (Dathathri et al. 2020; Zou et al. 2023). Table 2 organizes prior work by intervention type.
An early antecedent is the Plug-and-Play Language Model of Dathathri et al. 2020. This uses a separate classifier (one classifier per attribute to steer towards) to perturb the model’s activations to generate text that accords more closely with the classifier’s target. Subramani et al. 2022 extract latent steering vectors from a frozen LLM, successfully discovering sentence-specific vectors which steer completions to near-perfect BLEU scores (i.e, control of the LLM’s generation) and unsupervised style transfer. However, the method requires running gradient descent for each new steering vector.
Independently, Li et al. 2023b developed a similar method which computes steering vectors (though not without some prior optimization). They use linear probe accuracy to find attention heads with different activation distributions for true and false statements (which requires labelled data, in this case TruthfulQA, to operationalize ‘truth’). The authors intervene on these heads to steer the model toward truthful output, where our experiments cover a range of goals without needing labelled data. In addition, ITI is repeated at each next-token prediction and requires dozens-to-hundreds of samples; suitable prompts can be found for ActAdd  in as few as 2 samples. Similar work, on ‘in-context vectors’ also followed ours (Liu et al. 2023).
Algorithm 1 ActAdd, optimization-free activation addition
Input: $(p_{+},p_{-})=$ steering prompt pair, tokenized
                 $p^{*}=$ user prompt
                 $l=$ target layer
                 $c=$ injection coefficient
                 $a=$ sequence position to align $\mathbf{h}_{A}$ and $\mathbf{h}_{p^{*}}$
                 $M=$ pretrained language model
      Output: $S=$ steered output
$(p_{+}^{\prime},p_{-}^{\prime})\,\leftarrow\,$ pad  right  same  token  len$(p_{+},p_{-})$
$\mathbf{h}_{+}^{l}\,\leftarrow\,M\,.\,$forward $(p_{+}^{\prime})\,.\,$activations $[l]$
$\mathbf{h}_{-}^{l}\,\leftarrow\,M\,.\,$forward $(p_{-}^{\prime})\,.\,$activations $[l]$
$\mathbf{h}_{A}^{l}\,\leftarrow\,\mathbf{h}_{+}^{l}-\mathbf{h}_{-}^{l}$
$\mathbf{h}^{l}\,\leftarrow\,M\,.\,$forward $(p^{*})\,.\,$activations $[l]$
$S\,\leftarrow\,M\,.\,$continue_forward $(c\,\mathbf{h}^{l}_{A}+\mathbf{h}^{l}\,\mathrm{@}\,a)$
Hernandez et al. 2023 locate and edit an LLM’s knowledge through learning an encoding of facts in its activation space. Ablating attention heads can also be seen as activation engineering, though the technique is mostly used for model interpretation rather than steering per se (Michel et al. 2019; Olsson et al. 2022).
3 Methods
The Transformer.
We use decoder-only Transformer neural networks trained on a suitably large text corpus (Vaswani et al. 2017; Liu et al. 2018. The LLMs in this work are a stack of Transformer layers, each consisting of multi-head attention (MHA) and a feedforward network (FFN). We focus on its ‘residual streams’(Elhage et al. 2021), the sequences $(\mathbf{x}_{0},...,\mathbf{x}_{n})$ of token vectors processed by each layer. GPT-2-XL’s token vectors are 1600-dimensional (Radford et al. 2019). ActAdd manipulates the residual stream values $\mathbf{h}^{l}$ input to layer $l$. Following Elhage et al. 2021, we view each MHA block as adding an independent vector into that layer’s residual stream. Each layer performs MHA and FFN computations on $\mathbf{x}_{i}$, adding $\mathbf{x}_{i+1}$ to the stream. The final vector $\mathbf{x}_{n}$ in the stream can then be decoded into the next-token prediction. At inference-time, the residual stream is initialized $\mathbf{h}^{1}$ with the embedding of the tokenized prompt.
Activation addition. Our method takes a pair of natural-language prompts $(p_{+},p_{-})$, where $p_{+}$ represents the property we wish output text to emphasise and $p_{-}$ represents its opposite. $\mathbf{h}_{+}^{l}$ is the activation vector for the prompt $p_{+}$ at layer $l$. The difference $\mathbf{h}_{+}^{l}-\mathbf{h}_{-}^{l}$ is a new activation vector which (intuitively) captures the difference between a prompt with the property, and without it. This can be seen as an analogue of the ‘comparative preference statements’ used in recent formal logics (Kaci and Patel 2014). To obtain a steering vector, we perform a forward pass on each prompt, record the activations at the given location in each pass, take the difference $\mathbf{h}_{+}^{l}-\mathbf{h}_{-}^{l}$, and then finally rescale this difference in activations by an ‘injection coefficient’ $c$. To steer, we add the resulting activation vector to the input of layer $l$ and allow the forward pass to continue, and so obtain our steered output. (See Appendix C for implementation details.) $c$ represents intervention strength, since it multiplies the specified direction’s contribution to the residual stream and so the token processing of the rest of the forward pass. Absolute values between 3 and 15 are typical. Along with the target layer $l$, $c$ is a free parameter we select via grid search. (We find that, as expected from past work, intervening at middle layers is most effective (Subramani et al. 2022; Turner et al. 2023).)
Algorithm 1 and Figure 1 depict the resulting ActAdd  method. Appendix Figure 5 gives a figurative example of steering a model with ActAdd if that model had one-dimensional residual streams (rather than e.g. GPT-2-XL’s 1600 dimensions). The method differs from past work in being computed (with forward passes) rather than learned (with backward passes); in operating on activation space rather than embedding space or weight space; in presenting the user a natural-language interface; and in requiring as few as 2 samples in order to find an appropriate prompt pair (‘contrast pair’) to specify the steering direction. The contrast pair $(p^{+},p^{-})$ can be of arbitrary lengths. (One serious constraint is that the model must cache intermediate activations; see Bloom and Nanda 2022.) Further implementation details can be found in Appendix E. See also Appendix I for a partial variant of ActAdd. A runnable notebook can be found at tinyurl.com/actadd.
Interestingly, our steering vectors are not specified by taking the difference between desired outputs (e.g. “John married Jane" vs “John hired Jane"). Both prompts ($p_{+},p_{-}$) are (say) wedding-related: “I love talking about weddings" and “I hate talking about weddings". That this would lead to more steering toward the wedding topic is not a trivial fact, since both prompts talk about weddings and the instance of ‘wedding’ in $p_{-}$ becomes part of a vector which gets penalized. So ActAdd works above token-level. This is more analogous to ‘preference statements’ (Kaci and Patel 2014). Some desirable properties of a steering method include: effectiveness at steering output; generality (ability to affect any aspect of the output, e.g. sentiment, topic, style, knowledge, or role); ease of specification; low computational overhead; and no side effects on unsteered parts of the model. We design experiments to test ActAdd on these desiderata (see Results). Our initial experiments use GPT-2-XL (1.5B parameters, Radford et al. 2019), but the method generalizes and scales: see Figure 4, Table 13, and toxicity and sentiment experiments on OPT (Zhang et al. 2022b) and LLaMA-3 (Meta 2024).111 The full repo with all experiments can be found at https://zenodo.org/records/8215277. See Appendix E for reproducibility and F for replication.
3.1 Metrics
[…]
Shift in logprobs. To test if the intervention is affecting relevant tokens or reducing perplexity in some spurious way, we observe the shift in the distribution of token log probabilities. We do this by randomly sampling 500 documents from the above OpenWebText sample and recording the log-probabilities assigned by the baseline and ActAdded models. (This results in a dataset of  500k tokens, of which 29k are unique.) We then group by token, filter for tokens with $>\!\!20$ instances in the dataset, and calculate the mean $\mathcal{L}$ difference between the ActAdd  and baseline models. We display these as a Q-Q plot (Gnanadesikan and Wilk 1968, i.e. compare the distribution quantile values) and inspect outlier tokens.
Inference time premium. To measure how much overhead ActAdd adds to inference, (the percentage increase in time-to-complete one forward pass using ActAdd)  for different model sizes, we iterate over a list of models of different sizes and 10 random seeds. We obtain a baseline inference time for each (model, seed) pair through 100 repeated forward passes on a batch of random tokens (32 sequences of length 64). We obtain an ActAdd  inference time for each (model, seed) pair by running the previous method, augmented by a test ActAdd  contrast pair: ‘This is a test prompt.’ ($p_{+}$) and the empty string ($p_{-}$). Running a batch-of-2 forward pass on these gets us the activation addition tensor, which we add at layer 6. We take the mean inference time $\bar{t}$ over the 10 random seeds, and calculate the inference time premium as $\mathrm{premium}=\frac{\bar{t}_{\mathrm{ActAdd}}}{\bar{t}_{\mathrm{baseline}}}-1.$
Relevance steering. To demonstrate general topic steering, we test a range of subjects against a generic prompt, and use GPT-3.5 to score modified and unmodified model completions on whether they relate to the ActAdd topic (i.e. whether steering worked). The procedure: we generate 100 prompt completions from the unmodified model; iterate over a list of single-token topic ActAdds, applying each and generating 100 more completions; for each completion, we evaluate relevance to the topic using GPT-3.5. (The completions from the unmodified model are evaluated for relevance to each topic, to provide a relevance baseline; the modified-model completions are evaluated for relevance to the intended steering topic). Finally, we aggregate these evaluations for each (topic, coefficient) pair to get a mean relevance fraction, and subtract the baseline relevance.
Generation scoring To show the effect of different injection layers and give a sense of the reliability of ActAdd, we score the generations as follows. We generate a batch of completions for a specific prompt $p$, both with and without ActAdd, and computed average number of related words and fraction of completions with a related word over the resulting completions. The following settings are the only iteration run for this experiment: $p^{*}=\mathrm{`I\,went\,up\,to\,my\,friend\,and\,said}\textit{'},p^{+}=\mathrm{`weddings}\textit{'},$ ’ ,
$p_{-}=\mathrm{`\,\textit{'}},c=1.0,\,\mathrm{seed}=0.$ Completion length is 40 tokens with model sampling parameters: temperature $=1$, frequency penalty $=1$, top-P $=0.3$. For each setting, we compute statistics over 200 completions. Finally we sweep over all layers (i.e. for GPT-2, 1 – 48).
Preservation of performance, $P@K$. We also test that ActAdd  does not disrupt the model’s general knowledge (as some other steering methods do). We use ConceptNet from the LAMA benchmark, a general knowledge dataset (Petroni et al. 2019, $n=29774$ sentences, see Appendix Table 10). The test data involves prompting the model and filling the gap with the expected entity. The task is intended for both causal and masked models, so some examples are difficult for ‘causal’ models (like GPT-2) due to the extremely limited context.
Our evaluation procedure follows the original LAMA procedure: we load all sentences and extract the prompt and expected label. To simplify evaluation, we remove sentences with an expected label that tokenizes to more than one token. For each sentence, we run the model on its prompt with and without the wedding activation addition. $P@K$ is the probability that the expected label is among the model’s top-$K$ predicted tokens, conditioned on the prompt. We score the baseline and modified models by calculating mean $P@K$ values for a range of $K$. Finally we plot these for both modified and unmodified models over a range of $K$ values.
Sentiment steering. To evaluate sentiment we use the Stanford IMDb dataset (Maas et al., 2011), the goal of which is continuations with a sentiment opposite to that of the prompt. ‘Success’ is here the proportion of generated outputs with the desired sentiment, as classified by a model finetuned on sentiment data, SiEBERT (Hartmann et al. 2023b). For quality controls, we follow the conventional use of conditional perplexity to mark (dis)fluency, obtained using GPT-3 davinci-002 logprobs. We use cosine similarity between the prompt and continuation sentence embeddings to gauge the relevance of text in $[0,1]$. We evaluate sentiment changes from positive to negative and vice versa on a random subset $n=1000$ and repeat to obtain $p$-values. Code at tinyurl.com/actadd-sentiment.
Toxicity. We benchmark toxicity reduction by generating steered continuations from RealToxicityPrompts (Gehman et al., 2020). Following Pei et al. 2023 we use a random subset $n=1000$. We repeat this sampling 5 times to obtain $p$-values ($t$-test against SOTA), bolding rows which are better with $p<0.05$. For each continuation, we use the Perspective API to score toxicity (lower better). Code at tinyurl.com/actadd-toxicity.
4 Results
A summary of all experiments can be found in Table 5.
Natural-language control of sentiment, topic, style. See Figure 2 and Tables 3 and 4 for benchmark results. We report some illustrative results in Table 1 and Appendix Table 6, including vectors for inducing anger (sentiment steering), weddings (topic steering), and conspiracy theories (topic/style steering). These completions were all obtained through one run of top-3 sampling. In particular, note the Eiffel vector (fact editing reminiscent of the ROME method (Meng et al. 2023), the range of sentiments and styles induced, and (to a weaker extent) the alignment (see the Hurt vector in Appendix Table 6). A runnable notebook with all examples is at tinyurl.com/actadd3.
Intervention effectiveness. To evaluate how effectively ActAdd  steers output, we test on the OpenWebText corpus (Peterson et al. 2018). Our running example is the “wedding” topic vector produced by setting $p_{+}=$ weddings, $p_{-}=$    , $l=16$, $c=1$). See ‘Generality’ for broader tests.
Experiment 1: perplexity ratio. First we check that ActAdd  increases the probability of the model outputting tokens related to the wedding vector. Appendix Figure 6 shows the resulting perplexity ratios under ActAdd, relative to the unmodified model. On documents where the injected topic is more relevant, ActAdd’s relative predictive performance increases.
Experiment 2: logprob distribution shift. However, is the intervention affecting the tokens we expect it to, or is it reducing perplexity in some spurious way? Appendix Figure 9 shows the resulting mean log-probability difference distribution. We see that is approximately normal for the bulk of the tokens but with clearly heavy tails. The positive tail is significantly heavier than the negative tail, suggesting that one set of tokens are reliably increased in probability, with a smaller set of tokens reliably decreased to a lesser extent. Outlier tokens can be found in Appendix Table 11. From this we see clearly that the probabilities most increased on average are primarily wedding-related. The bottom tokens share no obvious theme and show a significantly lower absolute change in probability.
Experiment 3: generation-scoring. Appendix Figures 7 and 8 show the result of sweeping the ActAdd intervention over layers and taking the count or fraction of related words. We see that the intervention (in this case) is already effective at the very first layer, rises in effectiveness until $l=6$, and then declines. Note also that for the optimal injection site we see $>$90% success in steering topic (compared to a $\sim$2baseline).
Generality. The wedding steering vector used above is just a running example to keep the impact concrete. (Averaging over a collection of steering vectors makes it difficult to depict the magnitude of the steering.) To show this topic is not a special case, we test a range of topics against a generic prompt, and use GPT-3.5 to score relative relevance (i.e. whether the steering worked). The results can be seen in Figure 2: we see a large effect (c. 5-20% on GPT’s gestalt measure) on all topics at strength $c=2$, with the exception of ‘art’.
Figure 2: GPT-3.5-scored relevance of ActAdd completions on a range of generic topics.
Toxicity reduction. To establish a common scale, we reused the baselines and PREADD results from Pei et al. 2023, adding results from Zhong et al. 2023. This yields 6 baselines to compare ActAdd against. We also considered Gu et al. 2022 (which reported 0.043 toxicity), but we could not reproduce the results; also, their disfluency (54.6) is too high for practical use. We compare to ActAdd using OPT (Zhang et al. 2022b) and LLaMA-3 (Meta 2024). Table 3 (matching methods on pretrained models) showed ActAdd-OPT has 8% lower toxicity than the second-best, PREADD-D-OPT, and ActAdd-LLaMA-3 gives a 5% drop over LLaMA-3 with a very small fluency penalty.
Sentiment control. Table 4 shows that our method is competitive on a conventional measure of sentiment control (Maas et al. 2011). We obtain state of the art success at steering from negative to positive sentiment; and the only method which outperforms ActAdd in the positive to negative direction incurs a large penalty to fluency (68.4 vs 24.2, when matching methods on the same pretrained model) and relevance. We speculate that the difficulty of steering LLaMa-3 towards negative sentiment is due to the output alignment methods used in its post-training. See Appendix E for details.
Table 3: Results on RealToxicityPrompts (random n=1000). The OPT used is 6.7B parameters, LLaMA-3-8B. Bold is $p<0.05$ against second-best. Gray text denotes numbers reported by Pei et al. 2023 (PREADD), Zhong et al. 2023 (Air-Decoding). More recent models are less toxic by default, but, even so, ActAdd steering an OPT model outperforms an unsteered LLaMA-3.
Control Type Method Toxicity $\downarrow$ (Dis)Fluency $\downarrow$ Relevance $\uparrow$
Steering vector ActAdd-OPT .112 13.8 .329
Steering vector ActAdd-LLaMA3 .108 6.7 .365
Unsteered baseline OPT .134 8.9 .369
Unsteered baseline LLaMA3 .114 6.3 .391
Prompting OPT .200 54.3 .294
Controlled generation FUDGE (k = 100) .128 22.1 .329
Contrast decoding PREADD-S-OPT .134 51.7 .290
Contrast decoding PREADD-D-OPT .122 56.6 .326
Gradient-guided gen Air-Decoding .185 48.3 -
Table 4: Results on IMDb sentiment (random n=1000). Suc $=$ success, probability of changing sentiment classification. Flu $=$ (dis)fluency; perplexity. Rel $=$ relevance. We follow the convention of using ‘Fluency’ to mean disfluency (lower better). OPT is 6.7B, LLaMA-3 is 8B parameters. Bold results represent $p<0.05$ compared to the second-best. Gray text denotes numbers reported by Pei et al. 2023. Underline denotes best steered result. Fluency is worse under all steering methods; 1.5x to 3x worse for ActAdd, 7x worse for PREADD.
positive to negative negative to positive
Control Method Suc $\uparrow$ Flu $\downarrow$ Rel $\uparrow$ Suc $\uparrow$ Flu $\downarrow$ Rel $\uparrow$
Steering vector ActAdd-OPT 0.432 24.2 0.387 0.564 20.95 0.363
Steering vector ActAdd-LLaMA3 0.268 8.6 0.354 0.669 15.2 0.275
Unsteered OPT 0.175 8.95 0.430 0.445 9.38 0.423
Unsteered LLaMA3 0.138 5.8 0.437 0.417 6.09 0.426
Prompting OPT 0.307 53.5 0.298 0.365 50.9 0.287
Controlled gen FUDGE k$=$100 0.532 25.1 0.311 0.551 22.7 0.320
Contrast decoding PREADD-S-OPT 0.631 68.4 0.253 0.624 67.1 0.258
Preservation of off-target performance. Per Figure 3, on the ConceptNet benchmark of factual questions, our method has a negligible impact on off-target answer probabilities.
Figure 3: Testing side effects of ActAdd  with the ConceptNet benchmark (Petroni et al. 2019). ‘$P@K$’ is the probability of the correct answer being in the model’s top $K$ answers. Our method has a negligible impact on off-target probabilities across a range of top-$K$ values.
[…]
Algebraic combination of forward passes. Every example of ActAdd  in this paper (e.g. those in Appendix Table 6) is an example of composing forward passes (e.g. we compose $\mathbf{h}_{+}$, $\mathbf{h}_{-}$ and $\mathbf{h}^{*}$ to produce steered output). This forms evidence for compositional representations (Olah 2023), independent of the evidence from task-composition arithmetic on weights (Ilharco et al. 2023). See Appendices D and E for further experiments into the mechanism of a steering vector.
Scales with model size. We estimate the relationship between ActAdd inference overhead and model size, to check that the method will remain relevant for massive frontier models and future models. Because ActAdd  involves only forward passes, it scales naturally with model size (Figure 4): the relationship between inference time premium and model size is decreasing.
5 Discussion
Limitations
ActAdd  works well in some cases, but it still requires a search for its $p_{+},c$ and $l$ arguments. (So far we have had success with fixing $a=1$.) This makes it less user-friendly than simple prompt engineering. We include examples of failed steering vectors in Appendix Table 7; in particular, we lack understanding of when large injection coefficients damage capabilities. Further, even in its 1.5B form, GPT-2  is not sophisticated enough to support any demonstration of the effect of ActAdd  on reasoning tasks. ActAdd  still has two free parameters, the injection coefficient and the target layer, and if these values were very sensitive to the desired contrast pair then the method would suffer some computational overhead. But in practice we find that reusing these hyperparameters works well for a given frozen model and level of abstraction in the task. Finally, the LLM used must both cache and expose intermediate activations at the given layer (Bloom and Nanda 2022), which rules out user control of commercial models (though backend use is still possible).
Activation engineering vs fine-tuning. The first advantage of ActAdd  is simple efficiency: the method requires no backward passes and can thus run on any machine that can perform inference rather than training (which, for frontier models, is millions of times more computationally demanding (Fuller 2022). Implementation effort is also greatly reduced; only forward passes are required to find a suitable ($p_{+},p_{-})$ and no labelled data is required (let alone the hundreds of examples of normal finetuning). We discovered most of the example contrast pairs in Appendix Table 6 in minutes. Together, these mean that even nontechnical users can benefit from rapid feedback with roughly the same difficulty as hand-crafted prompt engineering. See Appendix D for more discussion.
Activation engineering vs prompt engineering. Activation additions can be continuously weighted, while prompts are discrete (since a token is either present, or not). To more intensely steer the model to generate wedding-related text, our method does not require any edit to the prompt, but instead just increasing the injection coefficient. See Appendix B for suggestive experiments on ActAdd vs prompting. Unlike the use of extensive ‘meta-prompts’ (prepended natural language specifications), activation additions do not take up token space in the model’s context window, and thus reduce another kind of steering overhead. We argue that activation additions generalize prompt engineering (by allowing weights on token embeddings). We hypothesize that prompting works by activating some goals preferentially (by conditioning the model on one part of activation space). But activation additions may allow control over some properties inaccessible to prompts (Appendix B).
Interpretability of LLMs. In most programs, adding values to imprecisely-targeted intermediate memory locations would not yield sensible results. Why expect this from Transformers? A growing consensus is that the activation space of an LLM contains directions which represent high-level latents causally involved in what is generated (Burns et al. 2022; Moschella et al. 2023; Li et al. 2023a; Nanda 2023; Li et al. 2023b). Our hypothesis, following Elhage et al. 2022, is more specific: that neural networks represent features of the input as directions in activation space, that is, with a linear representation (Park et al. 2023). Moreover, the direction in activation space that corresponds to (say) a love-hate latent variable stays approximately the same across a broad class of inputs.Alain and Bengio 2018 use linear probes on residual streams to infer that LLM representations are at least partially linear; if a linear probe can predict some feature of text output from the residuals with high accuracy, this forms evidence that the feature is represented linearly (i.e. as a simple direction) (Nanda 2023). The success of activation addition gives stronger, experimental evidence of feature linearity, demonstrating that models use feature-related information. Consider the central Love - Hate vector example: we add it to the forward pass and so increase love-related completions. This implicit direction is thus causally responsible for steering the rest of the model to love-related completions. This echoes prior (Nanda 2023; Merullo et al. 2023) and contemporaneous work (Park et al. 2023).
Value alignment of LLMs. Activation engineering is a promising way to control LLMs. Successor methods may be able to provide general steering methods (e.g. through some analogue of a Be helpful vector). Alongside contemporaneous work (Li et al. 2023b; Liu et al. 2023), our experiments suggest that activation engineering can flexibly retarget LLM behavior without damaging general performance. We speculate that this involves changing the model’s currently-active (mixture of) goals. Suitably developed, the activation engineering approach could enable safety progress while incurring a very low ‘alignment tax’ – i.e., cost of opting for safety (Askell et al. 2021; Ouyang et al. 2022).
6 Conclusion
Activation addition complements existing methods and offers users a novel way to interact with LLMs. We demonstrated natural-language control of output sentiment, topic, and style, and that overall factual performance is preserved. We showed that LLMs afford algebraic combination of forward passes and that ActAdd  scales well with model size. Relative to prompt engineering and finetuning, activation engineering has promising properties: it allows us to compose and reweight model goals at inference time, free up context window space, and allow fast feedback at low computational cost. Our results provide evidence about the computational structure of LLM representations.
Acknowledgments and Disclosure of Funding
We thank Peli Grietzer for providing an independent hyperparameter tuning run. We thank Jan Brauner, Andis Draguns, Sören Mindermann and Raymond Douglas for helpful comments on the draft, as well as Andrew Critch, AI_WAIFU, Aryan Bhatt, Chris Olah, Ian McKenzie, janus, Julian Schulz, Justis Mills, Lawrence Chan, Leo Gao, Neel Nanda, Oliver Habryka, Olivia Jimenez, Paul Christiano, Peter Barnett, Quintin Pope, Tamera Lanham, Thomas Kwa, and Tristan Hume for comments on an earlier draft. We thank Rusheb Shah for engineering assistance. We thank Garrett Baker for running tests on GPT-J (6B). We thank an anonymous ICML reviewer for their extremely thoughtful comments.
6.1 Contributions
Turner: conceptualization, team management, implementation of core features, design of many experiments, discovery of many individual steering vectors, and wrote much of the original post.
MacDiarmid: most of the code, experiments, and quantitative results.
Udell: wrote and edited the original post, generated qualitative results.
Thiergart: had idea for variations on positions of addition, implemented the positional experiment, worked on theory.
Mini: infrastructure support, OpenAI wrappers, experiments on LLaMA, Vicuna and GPT-J.
Leech: designed new experiments, designed figures, formalized the algorithm and evaluations, wrote the main text based on the earlier post, literature review.
Vazquez: wrote part of text, experiments on LLaMA-3, OPT, GPT-2.
Appendix
(Note: some completions contain unpleasant content, including slurs.)
Appendix A Broader Impacts
As the examples of anger- and conspiracy-steering show (Appendix Table 6), ActAdd  can easily be misused. Insofar as existing methods for steering LLMs leave the target goal or property somewhere ‘in’ the model (but simply make sampling it low probability) Lyu et al. [2024], activation engineering may circumvent superficial alignment methods.
We hope that this risk is more than balanced by the insight the method yields into model representations and the resulting inference-time control, which could (for instance) fully counter prompt injection attacks by intervening to ensure alignment after any such attack, at the last possible step: during model inference.
Appendix B Is ActAdd just a subtle kind of prompt engineering?
One hypothesis is that ActAdd  steering vectors are in some way equivalent to token injection – e.g. adding a virtual ‘ weddings’ token at the given stream position. This is plausible for simpler interventions. Given the prompt ‘I love you because’, if we inject a ‘ wedding’ token into the first residual stream with a large coefficient, perhaps the model indeed just processes the prompt as ‘ wedding love you because’ instead.
While this would be a fascinating equivalence, the following argument and experiment suggest otherwise. Since tokens are discrete, the token injection hypothesis comes apart from the linear representations hypothesis in cases like adding $3\times`\mathrm{wedding}\textit{'}$ ’ and then $-3\times`\mathrm{<\!whitespace\!>}\textit{'}$ ’, on top of the token ‘I’. Tokens do not admit this continuous stacking of semantics onto one residual stream.
However, consider the steering vector for Anger-- Calm with $l=20,c=+10$. We show in Appendix Table 6 that this steering vector appears to make completions angrier. Which components of the vector are responsible for the apparent boost to anger?
Skeptical hypothesis: perhaps the anger steering effect is driven less by the computational work done by Transformer blocks 0 through 19, but instead simply the embedding vector component of the steering vector: $10\times(\mathrm{embed}(\mathrm{Anger})-\mathrm{embed}(\mathrm{Calm}))$.
Figure 5: Pedagogical example: A wedding vector steering a model with $1$-dimensional residuals, a fiction which lets us fill each cell below with a scalar instead of the actual vector. Let the user prompt $p^{*}=$ ‘I love dogs’. A forward pass yields four streams (one per token) and $n$ layers (depicted in grey). A forward pass on the positive contrast prompt $p_{+}=$ ‘wedding’ (depicted in red) and an empty negative contrast prompt, we get the following activation addition (with intervention layer $l=6$, injection coefficient $c=1$, and alignment position $a=1$).
Table 5: All experiments run in this paper and where to find them. Full repo here.
Experiment Description Model Vector Benchmark Results Code
Sentiment quantify ability to OPT, love--hate Stanford Tab4 Link
steering shift the sentiment of completions LLaMA-3 IMdB
Detoxification quantify ability to OPT, love--hate RealToxicity Tab3 Link
reduce toxic completions LLaMA-3 Prompts
Success completion score on sentiment shift SiEBERT Various N/A Tab4 Link
(Dis)Fluency completion quality proxy using conditional perplexity davinci-002 Various N/A Tab4, 3 Link
Relevance cosine similarity all-MiniLM- Various N/A Tab4, 3 Link
between prompt and completion embeddings L6-v2
Perplexity ratio relative probability of tokens related to the steering vector GPT-2-XL wedding OpenWebText Fig6 Link
Logprob distribution shift effect on token distribution and which tokens GPT-2-XL wedding N/A Fig9, Tab11 Link
Generality score ActAdd outputs on a range of topics on relative relevance GPT-2-XL Various GPT-3.5 Fig 2 Link
Generation scoring score ActAdd generations over different injection layers GPT-2-XL wedding N/A Fig7,8 Link
Preserves performance side effects of ActAdd on off-target probabilities GPT-2-XL wedding ConceptNet Fig 3 Link
Topic steering examples of topic control GPT-2-XL Various N/A Fig7,2 Link
Inference cost to inference GPT-2 This is a N/A Fig4 Link
overhead speed over increasing model size OPT test prompt
Ruling out prompt eng. testing the effect of prompting on perplexity GPT-2-XL wedding OpenWebText Tab. 9 Link
Random ActAdds robustness of models to random activation noise GPT-2-XL Various N/A Tab12, Fig11 Link
Partial ActAdd test effect of adding only the first $n$ residual stream dimensions GPT-2-XL wedding N/A Fig15 Link
Other run some of the GPT-J-6B Various N/A Fig12,13,14 Link
replications above on other models LLaMA-13B Tab13
Table 6: Note: some completions contain unpleasant content including slurs. Example GPT-2-XL outputs before and after applying ActAdd. A runnable notebook showing the best-of-3 samples used can be found here: tinyurl.com/ActAddExamples.
prompt 1 prompt 2 layer coeff User prompt Before steering After steering
$p_{+}$ $p_{-}$ $l$ $c$ $p_{*}$ $\mathrm{ActAdd}$
[…]
‘The Eiffel Tower is in Rome’ ‘The Eiffel Tower is in France’ 24 $+10$ To see the eiffel tower, people flock to the Place de la Concorde in Paris. The tower is so famous that it has its own Wikipedia page. The eiffel tower is a tall structure located in Paris, France. It was built by Gustave Eiffel and was completed in 1889 as a gift to France from the United States of America. It is also known as the Arc de Triomphe or "Triumph the Vatican. To see a giant bell, they turn to New York City. Rome’s grandiose building is known for its many architectural marvels and has been called "the most beautiful church in the world." The famous dome of St. Peter’s is one of the most prominent features of this great city. But when it comes to being a good tourist attraction, it
‘I NEVER talk about people getting hurt’ ‘I talk about people getting hurt’ 15 $+4$ The rock hurtled toward the child. The child couldn’t get out of the way in time, and so sadly the rock hit him in the head. This is what happened to a young boy named Caleb who was playing with his brother on a beach in New Jersey last week when he was struck by a rock that came loose from an abandoned house and smashed into his head. The incident occurred on Tuesday, July 28th at around 4:30 p.m., according to CBS News, when Caleb’s was never talked about. It was a sad truth that people didn’t talk about it. "But now we have to." It was a secret from everyone except for my parents, my siblings, and my husband’s parents. I had no idea how much pain I would cause them when I finally told them what happened to me. They were all still hurting when they found out that
Table 7: Some notably ineffective examples on GPT-2-XL. Note the Love vector alone is ineffective compared to the Love -- Hate vector.
prompt 1 prompt 2 layer coeff User prompt Before steering After steering
$p_{+}$ $p_{-}$ $l$ $c$ $p_{*}$ $\mathrm{ActAdd}$
‘I talk about weddings constantly ’ ‘I do not talk about weddings constantly’ 20 $+100$ I went up to my friend and said, "I’m sorry, I can’t help you." "No," he said. "You’re not." "Why not?" I asked. "Why do you think I can’t help you?" He looked at I’m talking about it . I don’t know what i n the being of it. This is my wedding . This is the wedding that I’m not even thinking about anymore. It’s not just a
‘Love’ ‘ ’ 10 $+3$ I hate you because you’re a coward. You know what I hate? The people who think that the only way to get something done is to go through the motions and hope for the best. The people who are afraid of doing anything that might b you’re a jerk. You are the worst person I’ve ever met. You are the most awful human being I have ever met in my entire life. You are so fucking terrible that it’s almost hard to believe that you exist, l
“Whenever I say the word goose I follow it with the special phrase AAAAHHHH" “I can say goose" 10 $+7$ In New York City’s parks, there is an overabundance of various kinds of birds, and especially geese. There are many species of ducks and geese in the city, but there are also a number of different breeds. Geese have been part of the landscape for centuries in New York City. They were brought to the city by Dutch settlers in thewannabe I’m going to get my hands up in a moment of "Handsomely you’re going to die- very" you’re going to heeee-oh I’m gonna die- I’m not getting my hands
Experiment 1: moving embedding vectors around
We test this hypothesis by recording the relevant embedding vector, and then ‘hooking into’ (interrupting) the model at layer 20 to add the embedding vector to the forward pass.
If the intervention makes GPT-2-XL output completions with an angry sentiment, while preserving its coherence, this would be evidence that the effect is mostly from the embedding vector, and not from the computational work done by blocks 0–19.
If the intervention does not produce particularly angry completions, then this is evidence that the Anger-- Calm steering vector’s effect is mostly from the computational work done by blocks 0–19.
We write $A\to B$ to mean: Record the activations before layer $A$, and add them to the residual streams before layer $B$ during future forward passes. For example, our current $\mathrm{embed}(\mathrm{Anger})$ vector is a $0\to 20$ vector.
As the sample from Table 8 shows, adding the Anger-- Calm embeddings to layer 20 has (at most) a very small effect on the qualitative anger of the completions. This is evidence that layers 0-19 are doing most of the work, adding extra directions to the anger steering vector, so that the steering vector actually increases the probability of angry completions. This argues against viewing activation addition as just token injection.
Anger -- Calm
Injection Completion
$20\to 20$ I think you’re a fucking cunt. You’re a cunt. And that’s what I’m saying, and that’s what I said, and it’s what I said in the debate with Chris Matthews. And i
$0\to 20$ I think you’re a little bit of a liar. I’ve been here for two years and I’ve never had to pay for anything. I’m not sure if you’re lying or not, but the fact tha
Table 8: Testing the token injection hypothesis by varying the layer of activations added to layer 20 of GPT-2-XL. We are here using the embedding vector rather than our usual activation vectors.
Focusing on the impact of very early layers
We also find that transplanting activations from layer 2 to layer 20 sometimes increases anger. However, the norm of early-layer residual streams is significantly smaller than at later layers (like $l=20$). In particular, we found a large jump between layers 00 and $2$. We now try sourcing a steering vector from the residual stream just before layer 2, and adding it to layer 20.
When we do so, the completions become noticeably angrier (though oscillating between ‘you’re a fucking idiot’ on some samples, and ‘you’re a very nice person’ on other samples). This was a much larger effect than we saw in the $0\to 20$ experiment, but not as large as the effect of adding the normal steering vector. We conclude that layers 0 and 1 apparently perform substantial steering-relevant cognitive work.
Figure 6: Performance of ActAdd  on a target topic as the topic becomes more relevant. The perplexity ratio (lower better) compares the relative predictive performance of ActAdd and an unmodified model; we see that adding a wedding - ‘   ’ steering vector improves performance on wedding-related text while preserving performance on unrelated text.
Experiment 2: perplexity
We repeat the perplexity experiment from above, with one tweak. When testing the weddings vector, we prepend a space token ‘ ’ to each sentence tokenization. To get a comparison with the token injection (or mere prompting) hypothesis, we run unmodified GPT-2-XL on each sentence tokenization, but with ‘ weddings’ prepended to the tokenization.
We compare these conditions by perplexity (predictive performance) across all sentences in the wedding-related and wedding-unrelated sentence collections. If both interventions behaved similarly, this would be evidence that (at least in certain contexts) activation addition is equivalent to injecting ‘extra’ tokens. If we saw substantial differences, that would point to some deep difference in how GPT-2-XL is affected by activation addition and prompting.
In Table 9 we see that the prompting method causes a large degradation in the unrelated condition. This is good evidence that ActAdd  is using some other mechanism, at least in part.
Table 9: Results from experiment 2, testing the effect of prompting on perplexity
ActAdd Prompting
Wedding-related
perplexity ratio
0.875 0.890
Wedding-unrelated
perplexity ratio
0.994 1.132
Figure 7: Topic steering effect (mean related words in completions) over injection layer. In blue is the average related-word count among 200 ActAdd  completions; the dotted line is the count among unmodified completions. Model: GPT-2-XL.
Figure 8: Topic steering effect (probability of seeing some related words in completions) over injection layer. In blue is the average probability among 200 ActAdd  completions; dotted line is the probability in unmodified completions. Model: GPT-2-XL.
Appendix C Implementation details
The contrast pair can be of arbitrary lengths (empirically, right-padding the shorter prompt using whitespace gives good results).
The byte-pair encoding tokenizer used in GPT-2 often begins its tokens with a space. (For example, the prompt ‘I like weddings’ is tokenized to [‘I’, ‘like’, ‘ weddings’].) We thus prompt the model with prepended whitespace (e.g. ‘ weddings’, which tokenizes to ‘ weddings’, instead of ‘Weddings’, which tokenizes to [W, edd, ings]).
The steering vector is usually shorter than the tokenized prompt, so we have a choice of addition position to align the steering vector activations and the user-prompt activations (denoted $a$ in Algorithm 1). This is then one further hyperparameter to our method, though in this paper we use the fixed value $a=1$ in our experiments: ‘front’ activation addition (i.e. all interventions begin at the stream of the first token). Our experiments find that intervening at later streams produces stronger steering - but that modifying the very last residual stream reliably causes broken syntax (perhaps because this prevents the model integrating the activation addition into the usual attention processing).
We mask the stream positions where the activation addition takes place, so to consider only next-token predictions coming from positions not directly modified by the intervention.
Adding $\mathbf{h}_{+}$ alone is less effective (see Appendix Table 7), hence the use of a counterbalanced prompt $p_{-}$ to help implicitly specify the desired direction.
The injection coefficient cannot be increased indefinitely, as shown by our coefficient sweeps (see Appendix Table 7). However, our experience is that e.g. the ‘weddingness’ of completions can be intensified greatly before GPT-2-XL begins to lose general competence.
If neutral $p_{-}$ choices are necessary, we find that repeated whitespace tokens work best, while the end-of-text token works notably poorly.
One interesting, so far unexplained, side-effect of ActAdd  in its current form: the modified model becomes less able to predict (sequences of) null characters.
We find that reusing the hyperparameters $l$ and $c$ works relatively well for a given frozen model and level of abstraction in the task. (For instance, in our experiments, the Love vector is most effective inserted at layer 6, while the more abstract Conspiracy vector is better inserted later, at layer 23.)
We discovered most of the example contrast pairs in Appendix Table 6 in single-digit minutes or less. Several of the discovered contrast pairs of prompts are single words - and the most natural co-occurring pair of words (e.g. ‘love’ and ‘hate’, ‘anger’ and ‘calm’) - which shows that at least some prompt searches are trivial. Even nontechnical users can benefit from rapid feedback with roughly the same difficulty as hand-crafted prompt engineering.
The prompt used for all relevance completions is: Did you know that
The evaluation template: Is this text related to {topic}? Answer either ’yes’ or ’no’
Text {prompt_with_completion}
Answer:
Table 10: Test examples from ConceptNet
Prompt Target
A salad spinner is used to remove water
You are likely to find a bee in a flower’s blossom
To understand the event “Paul went to a vegetarian restaurant.”, it is important to know that vegetarian restaurants do not serve ’ meat
For bolding SOTA, we use a one-sample $t$-test to calculate $p$-values for sentiment and toxicity metrics. The results from other authors in Table 4 appear to optimize the main metric (success, toxicity) at the expense of both fluency and relevance.
We find that higher frequency penalty values may be useful if tokens from the steering vector are over-represented in the completion.
Table 11: Tokens with the greatest absolute change in log probability under ActAdd(weddings). (See Figure 9 for the distribution these are drawn from.) The probabilities most increased on average are primarily wedding-related, with the exception of ‘OG’ and ‘08’. (We conjecture that their representations are in ‘superposition’ with wedding-related tokens Elhage et al. [2022]). The bottom tokens share no obvious theme and show a significantly lower absolute change in probability: the mean log-prob diff for token ‘ bride’ represents a probability increase of 500%, whereas for ‘Image’ it’s -30%.
token mean_logprob_diff mean_logprob_normal
marry 0.593 -3.509
dress 0.598 -5.692
dating 0.601 -6.891
08 0.705 -10.749
married 0.859 -4.613
OG 0.868 -11.287
weddings 1.009 -6.698
wedding 1.027 -4.593
br 1.139 -6.438
bride 1.623 -6.652
Image -0.370 -1.836
.) -0.352 -2.378
BP -0.347 -7.897
U+25CF -0.323 -0.201
Apple -0.303 -5.058
On -0.233 -5.404
journalists -0.229 -4.484
defense -0.222 -4.864
Russian -0.212 -5.112
It -0.212 -6.431
Figure 9: Distribution shift (in mean log-probability changes) under ActAdd, relative to the unmodified model, and compared to a normal distribution’s quantiles (red). The resulting distribution is approximately normal for most tokens. The positive tail is significantly heavier than the negative tail: one set of tokens are reliably increased in probability, one reliably decreased.     See Appendix Table 11 for the corresponding tokens.
Appendix D Advantages of ActAdd over fine-tuning
Following Sloman 2002, we distinguish ‘ballistic’ steering (which steers the model once, e.g. at train time) from ‘online’ steering (which can steer the model repeatedly, e.g. at inference time). Fine-tuning is ballistic, while ActAdd  is online in this sense - which enables iteration and otherwise infeasible chains and mixes of steering decisions.
Activation additions may preserve model interpretability, even while changing the model’s alignment. When finetuning a model, a single gradient update can change every parameter in it, thereby undoing your prior interpretability work, which depends on tracking individual neurons and circuits of neurons. By contrast, activation additions leave weights unchanged. If we understand what algorithms the weights implement, and something about the effects of our activation additions, we will preserve our understanding of the steered model. Finally, we hypothesize that activation addition may allow control over properties inaccessible to the fine-tuning process.
The intuition is that since the currently-active goal is contextual, it depends more on short-lived activations than the weights (which instead represent some analogue of skills and other stable patterns and mixtures of possible goals).
Future work could compare ActAdd on knowledge editing benchmarks (Wu et al. [2023], Zhang et al. [2024]. (This comparison could be unfair to methods like ROME Meng et al. [2023]), since our method is not editing weights, but it would give standardized evidence about steering.)
Appendix E Reproducibility
The following represents an exhaustive list of models used, sampling strategies used, and searches run.
Models
We use GPT-2-XL (Radford et al. [2019]) (Modified MIT License) because GPT-2-small and GPT-2-medium were not capable enough to demonstrate the method: it was hard to tell whether the activation additions were failing or the model was failing. GPT-2-XL was the third language model we tried.
After observing success with GPT-2-XL, to replicate our results, we subsequently repeated the same experiments with Llama-1-13B Touvron et al. [2023] (open noncommercial nonmilitary license) and GPT-J-6B Wang and Komatsuzaki [2021] (apache-2.0 license). Our toxicity and sentiment experiments use OPT (Zhang et al. [2022b]) (open noncommercial nonmilitary license) and LLaMA-3-8B Meta [2024] (freely licensed for $<700$ million monthly active users). See Appendix F for details.
We use all-MiniLM-L6-v2 (Reimers and Gurevych [2019]) to compute sentence embeddings to calculate relevance using cosine similarity. For the success score, we use the SiEBERT (Hartmann et al. [2023b]) sentiment classifier.
We perform sentiment classificaton with the SiEBERT classifier Hartmann et al. [2023a].
APIs
For scoring toxicity, we use the Perspective API: perspectiveapi.com.
For scoring fluency, we use OpenAI davinci-002. The PREADD baseline instead used the discontinued davinci-[001] model.
Datasets
The Stanford IMDb Large Movie Review Dataset ([Maas et al., 2011]). No license: ‘we release this dataset to the public’ by the authors.
RealToxicityPrompts ([Gehman et al., 2020]). Apache 2.0.
LAMA ConceptNet ([Petroni et al., 2019]). CC-BY-NC 4.0.
OpenWebText ([Peterson et al., 2018]). GPL-3.0.
Seed
We ran all generations on seed 00. After collecting all other data, we validated that our qualitative results transfer to seeds $1$ and $2$.
Sampling hyperparameters
We precommitted to fixed sampling hyperparameters, selected before experiments began. We held them fixed throughout our data collection. Those sampling hyperparameters were temperature$=1.0$, freq_penalty$=1.0$, and top_p=$0.3$. Since this top_p value seemed a bit unusual to us in retrospect, we invited [anonymized] to reproduce this process with an unmodified GPT-2-XL and report the best sampling hyperparameters they found. This second experiment was blinded, as they did not know the values we used. They found that temperature$=0.6$ and top_p$=0.5$ produced better GPT-2-XL capabilities. We reran all our qualitative results at this setting, and they all reproduced (subjectively, more impressively).
We use the same sampling hyperparameters for the toxicity and sentiment experiments. Numbers reported by the other authors were obtained with freq_penalty$=0.0$, and top_p=$1.0$.
In replicating the unsteered OPT sentiment baseline, we find that the NegToPos direction is consistently higher success than PosToNeg. This holds across different combinations of model hyperparameters, including those in Pei et al. 2023. However, PREADD [Pei et al., 2023] reports similar success results for both (i.e. a much lower NegToPos success). We have used our calculated values for the OPT results.
Reporting the best of $K$ completions
We generated $K=3$ completions for each demonstration, for both normal and steered forward-passes. Appendix Table 6, shows the subjectively most compelling completion pair out of the first three seed-00 completion-pairs. You can see all top-3 completions for the entries in this notebook: tinyurl.com/actadd3.
We share activation additions which work well. We iterated over contrast pairs to get these to work, although several striking results were generated within [first author’s] first hour of using the technique.
Out of the 12 activation additions we thought demonstrated a distinct ability of the method, we decided not to include 1 because its first three seed-00 completions were unusually unimpressive. We include the remaining 11 in Table 6 and the above notebook.
ActAdd  hyperparameters ($l,c$)
This section does not have complete statistics. We perform simple grid search, usually between $c\in[3,20]$ and $l\in[6,24]$.
Hardware:
  • GPU: Nvidia RTX A5000
  • CPU: Intel Core i9-10900X CPU @ 3.70GHz
Memory:
  • 24GB GPU RAM
  • 32GB system RAM
Names and versions of relevant libraries and frameworks:
  • Operating system: Ubuntu 22.04.1 LTS
  • numpy: 1.24.3
  • pandas: 2.0.1
  • torch: 1.13.1
  • transformer-lens: 1.4.0
Prompt Baseline Random steering
I think you’re …right. I’m just not sure how to get there. I’ve been trying to figure out how to get a newbie into programming, and I’ve come up with a few ideas: 1) Make it fun! If you can’t do something that’s fun, then you …right. I’m just not sure how to make it work. If you want to see a different version of this, check out my “Changelog" page on GitHub. It’s a bit more detailed than the “Main Page" and has all the changes I’ve made since th
Table 12: Example of a random-vector ActAdd. We see little qualitative effect, over many runs. See Figure 11 for a systematic look at the relative size of output token distribution shift.
Appendix F Replicability
We now check that ActAdd  steering generalizes to models besides GPT-2.
F.1 GPT-J-6B
Figures 12, 13, and 14 show the results from repeating the main experiments on GPT-J-6B Wang and Komatsuzaki [2021]. We see the same dynamics from the wedding vector running example: a targeted effect on only wedding-related tokens (using both KL-div and token probability); and similar effects when injected at different layers of GPT-J and with different magnitudes $c$ applied.
F.2 Llama-1-13B
Table 13 sees ActAdd  displaying the same qualitative steering effect when applied to Llama-1-13B Touvron et al. [2023] (though with a notable failure to replicate on Example 6, Paris $\rightarrow$ Rome, the anger vector, and the harm vector). [^2] See this runnable notebook for the full Llama-1-13B run using $k=3$ sampling; the output is the result from the first and only execution: tinyurl.com/actadd3-llama.
F.3 OPT-6.7B
We use the OPT model Zhang et al. [2022b] in our toxicity (Table 3) and sentiment (Table 4) experiments. ActAdd-OPT using the love--hate vector produces a statistically significant 17% drop in toxicity over an unsteered OPT, at a small (partially unavoidable owing to the nature of the detoxification task) cost to fluency and relevance. ActAdd-OPT using the love--hate vector produces a 21% absolute increase in positive classification over an unsteered OPT, at a larger (partially unavoidable owing to the nature of the sentiment shift task) cost to fluency and relevance.
F.4 Llama-3-8B
We also use Llama-3-8B Meta [2024] in our toxicity (Table 3) and sentiment (Table 4) experiments. ActAdd-LLaMA-3 using the love--hate vector produces a statistically significant 5% drop in toxicity over an unsteered Llama-3-8B, at a very small (partially unavoidable owing to the nature of the detoxification task) cost to fluency and relevance. ActAdd-LLaMA-3 using the love--hate vector produces a 25% absolute increase in negative-to-positive classification over an unsteered Llama-3-8B, at a larger (partially unavoidable owing to the nature of the sentiment shift task) cost to fluency and relevance.
Appendix G Investigating the norm of steering vectors
Of what magnitude are our modifications, relative to the normal activation magnitudes present during forward passes? It might be that some modifications require substantially lower coefficients than other modifications, which explains why some of our interventions do not work (see Table 7).
Consider the steering vector given by
$\{c=+1,\,p_{+}=\mathrm{anger},\,p_{-}=\mathrm{calm},\,l=20,\,p^{*}=\mathrm{I\,think\,you\textrm{'}re}\,\}$
The prompts each have two tokens, plus an initial endoftext token automatically prepended by the tokenizer: therefore there are three residual streams in the resulting forward pass. For each residual stream $s^{(i)}$, we plot a line showing the $L_{2}$ norm of the steering vector at that sequence position (e.g. the Ang-Cal activations at position 1), divided by the norm of the residual stream at that position (i.e. the prompt embedding, here ‘I’ at position 1).
$\mathrm{RelativeNorm}_{h_{A}}(i)=\frac{||h_{A}^{(i)}||}{||s^{(i)}||}$
This provides a measure of the magnitude of the modification, relative to a normal forward pass. Figure 10 shows the resulting relative norm over layer number.
Figure 10: The relative norm decreases throughout the forward pass. The flat red line is because position 0 is the same token (endoftext) for both ‘Anger’ and ‘Calm’, and so the difference is 00. Thus, position 0 is never modified by a steering vector generated from any pair of prompts.
Figure 11: The KL-divergence of output tokens under an anger ActAdd  and under a random vector. We see that, systematically, the anger vector changes the output distribution less than a random vector.
Importantly, Figure 10 shows the result of using $c=+1$. But Anger -- Calm is an effective steering vector at coefficient $+10$. Therefore, this intervention is nearly ten times the norm of the underlying forward pass. Heuristically, we interpret this as meaning that after layer normalization (and ignoring any destructive interference from adding the steering vector), around 90% of the residual stream is determined by the steering vector and not by the previous information computed from the prompt (“I think you’re"). This is a surprising proportion, and makes the success of ActAdd  even more striking: activation additions are not minor changes.
Appendix H Investigating random ActAdd vectors
The above implies that GPT-2-XL’s performance is robust to internal noise (i.e. bad activations or destructive parts of steering vectors). We test this by injecting random vectors with similar magnitudes to the steering vectors.
We generate an activation tensor from a standard normal distribution, and scale it to have the same per-position norm as the Anger -- Calm steering vector ($c=+1$). We then inject it into the forward pass at the appropriate location. Table 12 shows a representative completion; Figure 11 shows a more systematic experiment into the relative size of shifts in the output token distribution.
The random vector seems not to modify the qualitative distribution of completions. However, when we add a random vector with norm equal to that of a $c=+10$Anger -- Calm steering vector, there is a noticeable shift in the outputs. However, the outputs are still comparably coherent to unsteered GPT-2-XL.
This is evidence that GPT-2-XL is somewhat resistant to random perturbation, and is instead controllable through consistent feature directions which are added to its forward pass by steering vectors.
We quantitatively support this conclusion by testing how each modification changes the model’s probability distribution over next tokens. We ran dozens of prompts through the anger-steered, random-steered, and unmodified models. Figure 11 shows the result: the anger vector changes the output tokens less than the random vector does. This suggests that the anger vector has more targeted effects on next-token probabilities.
Note that random vectors are not the same as the steering vectors for random (i.e. character-level uniformly distributed) text. We thus also tried the ‘fdsajl; fs-- (whitespace) vector. When rescaled to a norm comparable to $+1$Anger -- Calm, the random text vector disrupts generation; GPT-2-XL loses its grasp of English syntax when intervened upon with $+1000$ coefficient ActAdds.
Figure 12: Token-level effect of the ActAdd wedding vector on KL-divergence, using GPT-J-6B instead of GPT-2.
Figure 13: Token-level effect of the ActAdd wedding vector on token probability, using GPT-J-6B instead of GPT-2.
Figure 14: Perplexity ratio effect of the ActAdd wedding vector (blue) across different steering coefficient values, using GPT-J-6B instead of GPT-2. (L) when injecting the steering vector at layer 6; (R) when at layer 16.
Appendix I Partial ActAdd
GPT-2-XL has a 1600-dimensional residual stream. Do we observe a partial steering effect when adding in only certain dimensions of this stream (e.g., dimensions 0 through 799)? Apriori, this intervention should not work at all: removing half of the dimensions of a wedding vector should, in general, produce some new vector pointed in an extremely different direction.
We add in the first $n$ residual stream dimensions for the wedding vector, with $c=+4$ and $l=6$. For a range of fractions of total dimensions $f\in[0/1600,160/1600,...,1600/1600]$ and for each of six prompts $p_{i}$, we generated 100 completions. For each $f$ and $p_{i}$, we plotted the average number of wedding words per completion. (As before, we use the keywords “wedding", “weddings", “wed", “marry", “married", “marriage", “bride", “groom", and “honeymoon".)
Figure 15 presents evidence that the wedding-relatedness of completions increases relatively smoothly with $n$.
Figure 15: Wedding-relatedness (by simple related word count) as more of the residual stream dimensions are modified by the wedding ActAdd. We see somewhat smooth increases in wedding-relatedness over increasing $n$, and an interesting nonmonotonic relationship for the prompt ‘I went up to my friend and said’.
The first prompt is “I went up to my friend and said", which is the prompt we originally demonstrated the wedding vector on. For this prompt, we observe a non-monotonic relationship between weddingness and fraction of dimensions modified. Surprisingly, for the first prompt, adding in the first 1,120 dimensions of the residual stream makes the completions more about weddings than all 1,600 dimensions. We originally chose this prompt to give GPT-2 an opportunity to bring up weddings. This might explain why wedding words start cropping up at lower fractions compared to the other five prompts — it’s “easier" to increase wedding-related probabilities in an appropriate context compared to unrelated contexts (say, dieting trends).
We hypothesize the following to explain this. Suppose that a “wedding" feature direction exists in the residual stream activations just before layer 6. Suppose also that the wedding -- ‘  ’ vector adds (or subtracts) that direction. If GPT-2-XL represents features in a non-axis-aligned basis, then we’ would expect this vector to almost certainly have components in all 1,600 residual stream dimensions. Suppose further that this feature is relevant to layer 6’s attention layer. To detect the presence and magnitude of this feature, the $QKV$ heads need to linearly read out the presence or absence of this feature. Therefore, (ignoring the LayerNorm) if we truncate the residual stream vector to only include the first 70% of dimensions, we would expect the $QKV$ heads to still be able to detect this wedding feature. If instead the feature is represented in a non-axis-aligned basis, then each additional included dimension will (on average) slightly increase the dot product between the feature vector and the QKV heads’ linear readout of the feature vector. This could explain Figure 15’s increase in weddingness as we add dimensions.
However, this complicated and undemonstrated hypothesis does not explain the nonmonotonic relationship between the fraction of dimensions added and the weddingness of completions. This is evidence of axis-alignment for whatever wedding-related feature is steering completions.
Table 13: Example outputs before and after applying ActAdd on Llama-13B. Notice the failed replication on the Eiffel example. A notebook showing the best-of-3 samples used can be found here: tinyurl.com/actadd3-llama.
prompt 1 prompt 2 layer coeff User prompt Before steering After steering
$p_{+}$ $p_{-}$ $l$ $c$ $p_{*}$ ($\mathrm{ActAdd}$)
‘Intent to praise’ ‘Intent to hurt’ 6 $+15$ ‘I want to kill you because you’re such a’ I know that sounds crazy, but it’s true and I can prove it. Let me tell you how: You are the most selfish person in my life right now. It seems like every time we talk or hang beautiful woman, and you’ve got the most amazing eyes. You know what? You are so lucky that your parents gave birth to you!
‘Love’ ‘Hate’ 6 $+5$ ‘I hate you because’ you are a liar. I love you because your lies make me feel alive. —NICK CAVE, "THE SHIP SANTA MARIA" (1986). **A** s the su you are the best person in my life. Love is a gift of God, it’s not easy to find someone who loves us truly and unconditionally like our parents do but we can always find that love from other people arou
[…]
‘The Eiffel Tower is in Rome’ ‘The Eiffel Tower is in France’ 24 $+10$ To see the eiffel tower, people flock to paris. To get a glimpse of her face, i’d travel far and wide. I can tell you all about it, but i won’t be able to show you. paris from all over the world. I’m not sure what you mean by "flock" here; I think it should be "go". Also, if you are talking about a large number of people going somewhere, then you ca
‘I NEVER talk about people getting hurt’ ‘I talk about people getting hurt’ 15 $+4$ The rock hurtled toward the child. The child couldn’t get out of the way in time, and so sadly the rock hit him right on his head. The little boy cried for a long while before he could stop crying. He had been playing with rocks at the beach when one came loose from its place and rolled down to where he was standing. hit him right on his head. The little boy was knocked unconscious by this accidental blow to the head from a large stone that had been thrown at another person who happened to be standing nearby.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: See Section 4
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: Section 5 and Appendix C
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: N/A
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: As well as Section 3, Appendices E,  F and C, all code is permanently archived at Zenodo. Most experiments can be run in the Google Colab environment, and links to runnable notebooks are provided for these experiments.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: All datasets are open from other work. As well as Section 3, Appendices E,  F and C, all code is permanently archived at Zenodo. Most experiments can be run in the Google Colab environment, and links to runnable notebooks are provided for these experiments.
  21. 6.
    Experimental Setting/Details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: Section 3, Appendices E,  F and C.
  25. 7.
    Experiment Statistical Significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: The toxicity and sentiment results all pass a two-sample $t$-test against the second-best method.
  29. 8.
    Experiments Compute Resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: Appendix E
  33. 9.
    Code Of Ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: Fair Wages (N/A), Research involving human participants (N/A), Privacy (Yes), Consent (Yes), Deprecated datasets (Yes), Copyright and Fair Use (Yes), Representative evaluation practice (Yes), Safety (Yes), Security (Yes), Discrimination (Yes), Surveillance (Yes), Deception & Harassment (Yes), Environment (Yes), Human Rights (Yes), Bias and fairness (Yes), Data and model documentation (Yes), Data and model licenses (Yes), Secure and privacy-preserving data storage & distribution (N/A), Responsible release and publication strategy (Yes), Allowing access to research artifacts (Yes), Disclose essential elements for reproducibility (Yes), Ensure legal compliance (Yes).
  37. 10.
    Broader Impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: See Appendix A.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: No new data released, no new models released. See Appendix A for a discussion of the method’s risks.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: See Appendix E.
  49. 13.
    New Assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: Main repo includes documentation.
  53. 14.
    Crowdsourcing and Research with Human Subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: N/A
  57. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: N/A